OpenStudy (anonymous):

The measure of angle B is 65.8 degrees. Find the measure of the complement and of angle B. a.) 24.2; 155.8 b.) 155.8; 24.2 c.) 114.2; 24.2 d.) 24.2; 114.2

OpenStudy (anonymous):

complement of 90 = 65.8 - 90 24.2 complement of 180 = 180 - 65.8 = 114.2 Answer is D or C need to reword the question if you want an exact answer